OpenStudy (anonymous):

I need 3 irrational numbers between 4 and 5

zepdrix (zepdrix):

Using the square root is probably the easiest way to find irrationals.

zepdrix (zepdrix):

\(\Large\rm 4^2=16\) \(\Large\rm 5^2=25\) Therefore \(\Large\rm \sqrt{25}=5\) and \(\Large\rm \sqrt{16}=4\). (Because 16 and 25 are perfect squares). But what about the square root of the values between 16 and 25?

zepdrix (zepdrix):

The root of any value which is not a perfect square will be irrational. \(\Large\rm \sqrt{24}\) is irrational, and since it's between \(\Large\rm \sqrt{16}\) and \(\Large\rm \sqrt{25}\) that means it's between 4 and 5.